OpenStudy (anonymous):

Solve the following equation, explaining what you did in each step of your solution. After solving the equation, show the steps to check your solution. 3x + 2(4 + 6x) = 113

OpenStudy (anonymous):

Follow these steps:\[3x+2(4+6x)=113\]\[3x+8+12x=113\]\[15x+8=113\]\[15x=105\]\[x=7\]